# Semi-Supervised Fine-Tuning of Vision Foundation Models with Content-Style Decomposition

### v1

@@ Abstract
In this paper, we present a semi-supervised fine-tuning approach designed to improve the performance of foundation models on downstream tasks with limited labeled data. By leveraging content-style decomposition within an information-theoretic framework, our method enhances the latent representations of pre-trained vision foundation models, aligning them more effectively with specific task objectives and addressing the problem of distribution shift. We evaluate our approach on multiple datasets, including MNIST, its augmented variations (with yellow and white stripes), CIFAR-10, SVHN, and GalaxyMNIST. The experiments show improvements over purely supervised baselines, particularly in low-labeled data regimes, across both frozen and trainable backbones for the majority of the tested datasets.
@@ 2 Information-Theoretic Framework for Fine-Tuning of VFMs
We demonstrate the applicability of this approach by evaluating its performance across datasets that are simpler and distinct from large-scale, high-resolution natural image distributions, such as MNIST dataset variations[6], SVHN[8], CIFAR-10[4], and GalaxyMNIST[16]. More details about these datasets can be found in Section 2.2. Our results indicate that while certain vision foundation models struggle with these simpler datasets—particularly in low-labeled sample cases with purely supervised approaches—others are better suited to these specialized applications. This highlights the importance of model selection and fine-tuning strategies. This work showcases the advantages of our proposed approach in enhancing the adaptability and effectiveness of foundation models for handling diverse data.
@@ 3 Results and Discussion
We evaluated the performance of our proposed semi-supervised fine-tuning method across six datasets: MNIST, MNIST with yellow stripes, MNIST with white stripes, SVHN, CIFAR-10, and GalaxyMNIST. Experiments were conducted with both trainable and frozen backbones using three models: RADIOv2.5 Base [11], DINOv2 Small [9], and CLIP Base [3, 10, 1]. We focused on comparing our semi-supervised method to the purely supervised learning baseline, as it provides a well-established point of reference. Methods like LoRA[2] were left out of the scope of this work to avoid introducing additional parameters that would require extensive tuning, allowing us to prioritize a clearer evaluation of our semi-supervised fine-tuning approach. Results for each backbone are shown in Figures 3, 4, and 5.
@@ 3.1 Performance Across Datasets
Overall, our semi-supervised method consistently improves the performance across all datasets compared to the purely supervised approach. This is particularly evident in the low-labeled data regime (10-100 samples), where our method leverages the combination of labeled and unlabeled data to enhance the learning process. For simpler datasets like MNIST and CIFAR-10, frozen backbones benefit significantly from our method, while for more complex datasets like SVHN and GalaxyMNIST, the trainable models take advantage of the extra flexibility provided by fine-tuning.
@@ 3.2 Discussion: Frozen vs Trainable with Our Method
Overall, our method improves the scores compared to the purely supervised approach for most of the studied datasets with both frozen and trainable backbones.
@@ 4 Conclusion
Our experiments demonstrate that the proposed semi-supervised approach improves performance across both frozen and trainable backbones. The method consistently delivers better results than purely supervised baselines in the majority of the cases. Our method offers a step toward mitigating distribution shift, particularly in the early stages of fine-tuning. However, further research is required to fully understand how different backbone architectures respond to domain shifts and how unlabeled data can be leveraged more effectively in various scientific tasks.

### v2

@@ Abstract
In this paper, we present a semi-supervised fine-tuning approach designed to improve the performance of pre-trained foundation models on downstream tasks with limited labeled data. By leveraging content-style decomposition within an information-theoretic framework, our method enhances the latent representations of pre-trained vision foundation models, aligning them more effectively with specific task objectives and addressing the problem of distribution shift. We evaluate our approach on multiple datasets, including MNIST, its augmented variations (with yellow and white stripes), CIFAR-10, SVHN, and GalaxyMNIST. The experiments show improvements over supervised finetuning baseline of pre-trained models, particularly in low-labeled data regimes, across both frozen and trainable backbones for the majority of the tested datasets.
@@ 2 Information-Theoretic Framework for Fine-Tuning of VFMs
We demonstrate the applicability of this approach by evaluating its performance across datasets that are simpler and distinct from large-scale, high-resolution natural image distributions, such as MNIST dataset variations[6], SVHN[8], CIFAR-10[4], and GalaxyMNIST[16]. More details about these datasets can be found in Section 2.2. Our results indicate that while certain vision foundation models struggle with these simpler datasets—particularly in low-labeled sample cases with purely supervised fine-tuning approaches—others are better suited to these specialized applications. This highlights the importance of model selection and fine-tuning strategies. This work showcases the advantages of our proposed approach in enhancing the adaptability and effectiveness of foundation models for handling diverse data.
@@ 3 Results and Discussion
We evaluated the performance of our proposed semi-supervised fine-tuning method across six datasets: MNIST, MNIST with yellow stripes, MNIST with white stripes, SVHN, CIFAR-10, and GalaxyMNIST. Experiments were conducted with both trainable and frozen pre-trained backbones using three models: RADIOv2.5 Base [11], DINOv2 Small [9], and CLIP Base [3, 10, 1]. We focused on comparing our semi-supervised method to the purely supervised fine-tuning baseline, as it provides a well-established point of reference. Methods like LoRA[2] were left out of the scope of this work to avoid introducing additional parameters that would require extensive tuning, allowing us to prioritize a clearer evaluation of our semi-supervised fine-tuning approach. For all baselines in this paper (supervised or our method), we use pre-trained vision foundation models to ensure fair comparisons across methods. Results for each backbone are shown in Figures 3, 4, and 5.
@@ 3.1 Performance Across Datasets
Overall, our semi-supervised method consistently improves the performance across all datasets compared to the purely supervised fine-tuning approach. This is particularly evident in the low-labeled data regime (10-100 samples), where our method leverages the combination of labeled and unlabeled data to enhance the learning process. For simpler datasets like MNIST and CIFAR-10, frozen backbones benefit significantly from our method, while for more complex datasets like SVHN and GalaxyMNIST, the trainable models take advantage of the extra flexibility provided by fine-tuning.
@@ 3.2 Discussion: Frozen vs Trainable with Our Method
Overall, our method improves the scores compared to the purely supervised fine-tuning approach for most of the studied datasets with both frozen and trainable backbones.
@@ 4 Conclusion
Our experiments demonstrate that the proposed semi-supervised approach improves performance across both frozen and trainable backbones. The method consistently delivers better results than purely supervised fine-tuning baselines in the majority of the cases. Our method offers a step toward mitigating distribution shift, particularly in the early stages of fine-tuning. However, further research is required to fully understand how different backbone architectures respond to domain shifts and how unlabeled data can be leveraged more effectively in various scientific tasks.